An aqueous solution of sodium phosphate is a 0.142 M Na3PO4(aq) has a density of 1.02 g/ml What is the percent by mass of sodium phosphate ? What is the mole fraction of sodium phosphate
1
Expert's answer
2021-08-09T07:47:57-0400

The percent mass of Na3PO4 in this solution we can determine, using following equation:

w = C×M(Na3PO4) / (10×p), where C is molar concentration;

M is molar mass of sodium phosphate; p is density.

Thus, w = 0.142×(3×23 + 31×1 + 4×16) / (10 × 1.02) = 2.28 (%)

The equation that describes the connection between mole fraction and percent mass is given bellow:

x = (w/M(Na3PO4)) / ((w/M(Na3PO4) + ((1-w)/M(H2O))

Using given values,

X = (0.0228/164) / ((0.0228/164) + ((1-0.0228)/18) = 0.0026

Answer: w = 2.28%; X = 0.0026.